OpenStudy (anonymous):

limit as x approaches zero is sin^2x/x... find the limit of the trigonometric function

OpenStudy (anonymous):

Well, since plugging in 0 gets you the indeterminate form 0/0 you can use l'hopital's rule to find the limit

OpenStudy (anonymous):

what is that

zepdrix (zepdrix):

You haven't been taught L'Hopital's Rule yet? Hmm darn, I guess we're suppose to use Squeeze Theorem or something else here. Hmm.

zepdrix (zepdrix):

Hmm yah i think that works :)

zepdrix (zepdrix):

\[\Large \lim_{x\to0}\frac{\sin^2x}{x} \qquad=\qquad \lim_{x\to0}\frac{\sin x}{x}\cdot \lim_{x\to0}\sin x\]That first limit on the right is one you want to remember! :O

OpenStudy (anonymous):

Welp, deleted that by mistake, what I said was change to (sinx/x) * sinx and solve (knowing that sinx/x = 1)

OpenStudy (anonymous):

i get that sine of x over x is one but then one happens to the sine of x we took out

OpenStudy (anonymous):

Plug in 0 for x and you get your limit :)

OpenStudy (anonymous):

so would the answer be 3pi/2